Doreen Schmidt is a chemist with Gemco Pharmeceutical. She need o prepare 12 liters of a 9% hydrochloric solution. Find the amount of a 4% solution and the amount of a 12% solution she should mix to get this solution.
Answer by stanbon(75887)

She need o prepare 12 liters of a 9% hydrochloric solution. Find the amount of a 4% solution and the amount of a 12% solution she should mix to get this solution.
------
Equation:
acid + acid = acid
------
0.04x + 0.12(12-x) = 0.09*12
-----
Multiply thru by 100 to get:
4x + 12*12 - 12x = 9*12
----
-8x = -3*12
-----
x = 3(3/2) = 4.5 liters (amt. of 4% solution needed)
---
12-x = 7.5 liters (amt. of 12% solution needed)
=====================
Cheers,
Stan H.
================